I need help solving this problem.
10x^4-27x^2+5=0
Thanks

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

What sort of help?
Do you know about the quadratic equation?
Can you factor it?

Answer by stanbon(75887)

I need help solving this problem.
10x^4-27x^2+5=0
---
Factor:
10x^4 - 25x^2 - 2x^2 + 5 = 0
--------
5x^2[2x^2 - 5] -(2x^2 - 5] = 0
--------
Factor:
[2x^2-5][5x^2-1] = 0
x^2 = 5/2 or x^2 = 1/5
---
x = +-[sqrt(10)/2] or x = +-sqrt(5)/5
======================================
Cheers,
Stan H.
====================